# Generic Antiholomorphic Polymonial Vector Fields

### v1

@@ 3.1. Counting the number of noncrossing trees
Fix a number of vertices $n\geq 1$. Given a $n$-noncrossing tree, we index every vertex counterclockwise by 1 to $n$ (mod $n$), where $j$ (mod $n$) represents vertex $e^{(j-1)\frac{2\pi i}{n}}$. We define
@@
Our main interest is to compute a formula for $A(n)$. This is done more easily by introducing an intermediate quantity and to link $A(n)$ with it. We define
@@
To count $A_{1}$, remark that every nc tree that satisfies $\deg{1}=1$ must have one and only one edge $e$ coming out of $1$. Let $2\leq j\leq n$ be the other end of $e$. We get two subtrees $G_{1}$ with the $n_{1}=j-1$ vertices $2$ to $j$ and $G_{2}$ with the $n_{2}=n+1-j$ vertices $j$ to $n$. Each $G_{i}$ must be a noncrossing tree of order $n_{i}$ that has no edge that reaches $1$, so $G_{i}$ is a $n_{i}$-nc trees.
Conversely, for $2\leq j\leq n$, suppose we’re given two nc trees $G_{1}$ and $G_{2}$ of order $n_{1}$ and $n_{2}$ that satisfy $n_{1}+n_{2}=n$. Take the disk with $n$ evenly spaced vertices labelled from $1$ to $n$ on it’s boundary and draw the edge $(1,j)$. We can reconstruct a nc-tree of order $n$ by inserting $G_{1}$ above the edge $(1,j)$ and $G_{2}$ under $(1,j)$ (see Figure 5). The resulting nc tree must satisfy $\deg{1}=1$, hence is of the type we were looking for.
@@
Let $G\subseteq\overline{\mathbb{D}}$ be a finite embedded graph. For every connected component $X_{j}$ of $\overline{\mathbb{D}}\setminus G$, we define $e_{j}$ as the middle point of $X_{j}\cap S^{1}$ if the latter is non empty and as any point in $\mathring{X_{j}}$ if $X_{j}\cap S^{1}=\varnothing$. We say that two connected components are adjacent if their closure intersect in a whole edge of $G$.
@@
Let $G$ be a noncrossing tree. Using the same notation as above, we define it’s dual graph $G^{*}$ as the graph on the vertices $e_{j}$ and edges joining $e_{i_{1}}$ and $e_{i_{2}}$ if $X_{i_{1}}$ and $X_{i_{2}}$ are adjacent.
@@
The number of $G_{1}$ is $A(j-1)$. To count the number of $(n-j+2)$-nc tree $G_{2}$ that have $(1,j)$ as an edge, consider the dual tree $G_{2}^{*}$. Label its vertices by $j_{d}$ to $(n+1)_{d}$ counter clockwise, starting with the vertex between $1$ and $j$ (i.e. $\ell_{d}$ is at position $\exp\left({(\ell-j+1)\frac{2\pi i}{n_{2}}-\frac{\pi i}{n_{2}}}\right)$, where $n_{2}=n+2-j$). Then, to ask for $(1,j)$ to be an edge in $G_{2}$ is equivalent to ask $G_{2}^{*}$ to be a $(n+2-j)$-nc tree with $\deg(j_{d})=1$. Hence, there are $A_{1}(n+2-j)$ possibilities for the tree $G_{2}$. This gives Equation (3). ∎
Figure 6. There is a bijection between $n$-nc trees such that $(1,j)$ is an edge and $(1,\ell)$ is not an edge for all $12\leq\ell\leq j-$ 1 and the set of pairs $(G_{1},G_{2})$ where $G_{1}$ is a $(j-1)$-nc tree and $G_{2}$ is a $(n+2-j)$-nc tree that has $(1,j)$ as an edge. Furthermore, nc trees such as $G_{2}$ are characterized by having a dual (dashed) with $\deg j_{d}=1$.
@@
Then if $\mathcal{B}(z)$ is the generating function, the recurrence relation gives $z\mathcal{B}^{3}(z)=\mathcal{B}(z)-1$. Using Lagrange’s inversion formula then gives us that $B(n)=\frac{1}{2n+1}\binom{3n}{n}$ from which the proposition follows. This is done in [Noy98]. ∎
This number is well known, see for instance [Knu97], and correspond to the sequence A001764 up to a reindexing. Those numbers also count ternary trees, a fact that we will use in Section 6.1.
@@
We are now interested in $A^{r}(n)$, the number of $n$-nc-trees modulo rotations. We can obtain a formula for $A^{r}(n)$ in terms of $A_{1}(n)$, so we begin with $A_{1}(n)$.
Proposition 3.7.
The numbers $A_{1}(n)$ satisfy $A_{1}(1)=1$ and
$A_{1}(n)=\frac{2}{3n-4}\binom{3n-4}{n-2}.$ (5)
@@
This is done by expressing $A_{1}(n)$ in a convolution of some $A(j)$ and then using Lagrange’s inversion formula (see [Noy98]). ∎
Looking at some examples shows us that some nc-crossing tree may be obtained from others by rotation. The goal of the following discussion is to compute the number of nc trees modulo rotation.
Lemma 3.8.
Let $G$ be a $n$-nc tree, then $G$ has a non-trivial rotational symmetry $\alpha$ only if $n$ is even. Then, if this is the case, $\alpha$ necessarily fixes one and only one edge with antipodal endpoints.
Proof.
Suppose $\alpha$ acts on $G$ by rotation (i.e. it translates every vertex by a constant amount ($t$ mod $n$)). Index every vertex by 00 to $n-1$ (mod $n$). Let $m=\text{ord}\,\alpha$. Then we have $m|n$. We first claim that there exists some subgraph $H$ of $G$ such that $G=\bigcup_{0\leq j\leq m-1}\alpha^{j}H$. Indeed, let $d=\gcd(n,m)$ and let $S=\{0,...,d-1\}$. Define $H$ as the subgraph induced by the edges that have an endpoint in $S$. Then $H$ satisfy the claim.
@@
Now, suppose that $\alpha:\text{($j\mod n$)}\mapsto(j+\frac{n}{2}\mod n)$ is a rotational symmetry of $G$ and let $H$ be the subgraph as described above. Remark that there must be an external edge since, otherwise, we would have that #{edges of $\#\{\text{edges of $G$}\}=n-1=2i$, contradicting the parity of $n$.
Let $\pi:\mathbb{Z}\to\mathbb{Z}/n\mathbb{Z}$ be the projection and lift $G$ to $\mathbb{Z}$. Let $f$ be an external edge. Conjugating $G$ by a rotation, we may assume that $\tilde{f}:=(0,j)$ is a lift of $f$, with $\frac{n}{2}\leq j\leq n-1$. Recall that the rotation $\alpha$ lifts to a translation $\tilde{\alpha}$ by $\frac{n}{2}$. Consider the edge $\tilde{\alpha}.\tilde{f}=(\frac{n}{2},j+\frac{n}{2})$. Then, whenever $\frac{n}{2}<j\leq n-1$, the edge $\alpha.f=\pi(\tilde{\alpha}.\tilde{f})$ and $f$ cross in $G$, so we must have that $j=\frac{n}{2}$, meaning that $(0,\frac{n}{2})$ is an edge that fixes antipodal points. ∎
@@
$A^{r}(n)=\frac{1}{n}\bigg{(}A(n)+\mathbf{1}_{n\text{ even}}nA_{1}\bigg{(}\frac{n}{2}+1\bigg{)}\bigg{)}$ (7)
This is sequence A296532 in the OEIS. This result was probably already proved, but we could not find a reference. Therefore, we added a proof here. A discussion about similar cases may be found in [Noy98].
@@
In the case $n$ is odd, proposition 3.8 tells us that there are precisely $n$ elements in every rotation class.
When $n$ is even, every nc tree that has a rotational symmetry has $m:=\frac{n}{2}$ elements in it’s rotation class. So, we need to compute the number of nc trees that have a rotational symmetry to artificially add them up before dividing by $n$.
Index vertices of $G$ by $10\leq j\leq 2m-$ 1 (i.e. vertex $j$ corresponds to $e^{j\frac{2\pi i}{2m}}$).
By proposition 3.8, counting those amounts to first counting nc trees that have $f=(0,\frac{n}{2})$ as an edge and have a rotational symmetry.
Let $G$ be a nc tree on $2m$ vertices that is invariant under the order $2$ rotation $\alpha$ and $f=(0,m)$ the edge fixed by $\alpha$. Since $f$ splits $G$ in two $(m+1)$-nc trees $G_{1}$ and $G_{2}$ and that those two trees mutually determine the other one via $\alpha$, we’re left to compute the number of $(m+1)$-nc trees $G_{1}$ that have $(0,m)$ as an edge. By the same argument as in Lemma 3.4, going to the dual tree tells us that this number is $A_{1}(m+1)$. ∎
@@ 3.2. Characterization of the topological phase portrait with noncrossing trees
Given a polynomial $P\in\Xi_{k+1}$, the combinatorial invariant of $\overline{P}{\partial\partial z}$ is the $(k+2)$-nc tree induced by the separatrix graph.
@@
Let $P_{1},P_{2}\in\Xi_{k+1}$. If $\overline{P_{1}}{\partial\partial z}$ and $\overline{P_{2}}{\partial\partial z}$ have the same combinatorial invariant, then they are quasi-conformally equivalent. (In other words, there exists a quasi-conformal mapping $h$ that maps orbits of $P_{1}$ on orbits of $P_{2}$). In particular, the two vector fields are topologically orbitally equivalent.
@@
$\bigl{(}{t_{i}^{j}}_{*}\overline{P_{i}(z)}\frac{\partial}{\partial z}\bigr{)}_{t_{i}^{j}}=|P_{i}(z)|^{2}\frac{\partial}{\partial t_{i}^{j}},$
@@
Recall that the rectifying coordinate $t$ maps orbits to horizontal lines. Since the $\psi_{j}$’s we’ve just built map horizontal lines to horizontal lines, the composition $h_{j}:=41(t_{2}^{j})^{-1}\circ\psi_{j}\circ t_{1}^{j}$ maps orbits to orbits.
Denote by $\mu_{j}(z)=\frac{\overline{\partial}\psi_{j}(z)}{\partial\psi_{j}(z)}$ the Beltrami coefficient of $h_{j}$. One can compute that $\mu_{j}(z)=\frac{\tau_{2}-\tau_{1}}{\overline{\tau_{1}}-\tau_{2}}$ and the condition $|\mu_{j}(z)|<1$ is seen to be equivalent to $04\Im(\tau_{1})\Im(\tau_{2})>$ 0 which is true. Note that all $\psi_{j}$’s are invertible and extend to homeomorphisms between $\overline{A_{1}^{j}}$ and $\overline{A_{2}^{j}}$, where $A_{i}^{j}$ is either a strip or a half-plane. Hence, we have the following picture for every $j$.
@@
Now, the maps $h_{j}$ coincide along the closure of $X_{1}^{j}$ and $X_{1}^{j^{\prime}}$ (if ever) so we can glue them in a global homeomorphism $\Psi:\mathbb{C}\to\mathbb{C}$ that maps orbits to orbits. ∎
@@
Suppose that $\overline{P}_{j}\frac{\partial}{{\partial\partial z}}$ has combinatorial invariant $G_{j}$ for $j=1,2$. If there exists a rotation $\rho$ such that $\rho.G_{1}=G_{2}$, then $\rho_{*}\left(\overline{P}_{1}{\partial\partial z}\right)$ is topologically orbitally equivalent to $\overline{P}_{2}{\partial\partial z}$.

### v2

@@ 3.1. Counting the number of noncrossing trees
This section is devoted to counting the number of isomorphism classes of noncrossing trees of order $n$ rooted at $1$. Fix a number of vertices $n\geq 1$. Given a $n$-noncrossing tree, we index every vertex counterclockwise by 1 to $n$ (mod $n$), where $j$ (mod $n$) represents vertex $e^{(j-1)\frac{2\pi i}{n}}$. We define
@@
Our main interest is to compute a formula for $A(n)$. This is done more easily by introducing the intermediate quantity
@@
To count $A_{1}$, remark that every nc tree that satisfies $\deg{1}=1$ must have one and only one edge $e$ coming out of $1$. Let $2\leq j\leq n$ be the other end of $e$. Consider the two subgraphs $G_{1}$ with vertices $2$ to $j$ and $G_{2}$ with vertices $j$ to $n$ and set $n_{1}:=j-1$ and $n_{2}:=n+1-j$. There are no edges from $1\leq m<j$ to $j<\ell\leq n$ and no other edges to 1, so $G_{i}$ must have $n_{i}-1$ edges, therefore it is a $n_{i}$-nc tree.
Conversely, suppose we’re given two nc trees $G_{1}$ and $G_{2}$ of order $n_{1}$ and $n_{2}$ that satisfy $n_{1}+n_{2}=n$. Let $j:=n_{1}+1$. Take the disk with $n$ evenly spaced vertices labelled from $1$ to $n$ on it’s boundary and draw the edge $(1,j)$. We can reconstruct a nc tree of order $n$ by mapping the vertices of $G_{1}$ to the vertices $2$ to $j$ on $S^{1}$ and the vertices of $G_{2}$ to vertices $j$ to $n$ on $S^{1}$ (see Figure 5). The resulting graph is a nc tree that satisfies $\deg{1}=1$.
@@
The dual of nc tree $G$ is the usual dual of a planar graph, with the additional constraint that the vertices lie at position $e_{j}^{\prime}:=e^{-2i\pi 2k}e_{j}$ on $S^{1}$. We note it $G^{*}$.
For a $n$-nc tree $G$, there are $n$ connected components of $\overline{\mathbb{D}}\setminus G$, hence its dual $G^{*}$ has $n$ vertices. Every edge of $G^{*}$ must cross one and only one edge of $G$, therefore $G^{*}$ has $n-1$ edges, which implies that $G^{*}$ is a tree. Furthermore, $G^{*}$ is noncrossing since the dual of a planar graph is planar.
@@
The number of such $G_{1}$ is $A(j-1)$. To count the number of $(n-j+2)$-nc trees $G_{2}$ that have $(1,j)$ as an edge, consider the dual tree $G_{2}^{*}$. Label its vertices by $j^{*},(j+1)^{*},...,(n+1)^{*}$ counter clockwise, starting with the vertex between $1$ and $j$ (i.e. $\ell^{*}$ is the $(\ell-j)$-th vertex at position $\exp\left({(\ell-j+1)\frac{2\pi i}{n_{2}}-\frac{\pi i}{n_{2}}}\right)$, where $n_{2}=n+2-j$). Then, to ask for $(1,j)$ to be an edge in $G_{2}$ is equivalent to ask $G_{2}^{*}$ to be a $(n+2-j)$-nc tree with $\deg(j^{*})=1$. Hence, there are $A_{1}(n+2-j)$ possibilities for the tree $G_{2}$. This gives Equation (3). ∎
Figure 6. There is a bijection between $n$-nc trees such that $(1,j)$ is an edge and $(1,\ell)$ is not an edge for all $12\leq\ell\leq j-$ 1 and the set of pairs $(G_{1},G_{2})$ where $G_{1}$ is a $(j-1)$-nc tree and $G_{2}$ is a $(n+2-j)$-nc tree that has $(1,j)$ as an edge. Furthermore, nc trees such as $G_{2}$ are characterized by having a dual (dashed) with $\deg j^{*}=1$.
@@
Then if $\mathcal{B}(z)$ is the generating function, the recurrence relation gives $z\mathcal{B}^{3}(z)=\mathcal{B}(z)-1$. A consequence of Lagrange’s Inversion Formula states that a formal series $\mathcal{A}(z)=\sum_{n=0}^{\infty}a_{n}z^{n}$ satisfying $\mathcal{A}(z)=a_{0}+zg\bigl{(}\mathcal{A}(z)\bigr{)}$ will have coefficients
[…]
Applying this formula with $a_{0}=1$ and $g(z)=z^{3}$ to $\mathcal{B}(z)$ then gives us that $B(n)=\frac{1}{2n+1}\binom{3n}{n}$ from which the proposition follows. This is done in [Noy98].
The formula for $A_{1}(n)$ is done in a similar way; see [Noy98] for more details. The sequence $\bigl{(}A_{1}(n+1)\bigr{)}_{n\geq 1}$ is A006013 in the OEIS. ∎
The sequence $A(n)$ is well known and corresponds A001764 in the OEIS up to a reindexing. Those numbers also count ternary trees, a fact that we will use in Section 6.1. Indeed, let $T(n)$ be the number of ternary tree with $n$ internal vertices, since there are three subtrees attached to the root, it follows that $T(n)$ verifies Equation 6, and since $T(0)=1$, $T(1)=1$, we obtain $T(n)=A(n+1)$.
@@
Let $A^{r}(n)$ be the number of $n$-nc trees modulo rotations. Looking at some examples shows us that some nc-crossing tree may be obtained from others by rotation. The goal of the following discussion is to compute the number of nc trees modulo rotation.
Lemma 3.7.
Let $G$ be a $n$-nc tree, then $G$ has a non-trivial rotational symmetry $\alpha$ only if $n$ is even. Moreover, if this is the case, $\alpha$ necessarily fixes a unique edge with antipodal endpoints.
@@
Suppose $\alpha$ acts on $G$ by rotation (i.e. it translates every vertex by a constant amount mod $n$). Index every vertex by 00 to $n-1$ (mod $n$). Let $m=\text{ord}\,\alpha$. Then we have $m|n$. We first claim that there exists some subgraph $H$ of $G$ such that $G=\bigcup_{0\leq j\leq m-1}\alpha^{j}H$. Indeed, let $d=\tfrac{n}{m}$ and let $S=\{0,...,d-1\}$. Define $H$ as the subgraph induced by the edges that have an endpoint in $S$. Then $H$ satisfies the claim.
@@
Now, suppose that $\alpha:j\mapsto j+\frac{n}{2}$ (mod $n$) is a rotational symmetry of $G$ and let $H$ be the subgraph as above. Remark that there must exist an external edge $f$ since, otherwise, we would have that #{edges of $\#\{\text{edges of $G$}\}=n-1=2i$, contradicting the parity of $n$.
Let $\pi:\mathbb{Z}\to\mathbb{Z}/n\mathbb{Z}$ be the projection and lift $G$ to $\mathbb{Z}$. Conjugating $G$ by a rotation, we may assume that $\tilde{f}:=(0,j)$ is a lift of $f$, with $\frac{n}{2}\leq j\leq n-1$. The rotation $\alpha$ lifts to a translation $\tilde{\alpha}$ by $\frac{n}{2}$, so that $\tilde{\alpha}.\tilde{f}=(\frac{n}{2},j+\frac{n}{2})$. Now, whenever $\frac{n}{2}<j\leq n-1$, the edge $\alpha.f=\pi(\tilde{\alpha}.\tilde{f})$ and $f$ cross in $G$, so we must have that $j=\frac{n}{2}$, meaning that $f$ has antipodal endpoints. Since the graph is noncrossing, this edge must be unique. ∎
@@
$A^{r}(n)=\frac{1}{n}\bigg{(}A(n)+\mathbf{1}_{n\text{ even}}{n2}A_{1}\bigg{(}\frac{n}{2}+1\bigg{)}\bigg{)}$ (8)
This is sequence A296532 in the OEIS up to a reindexing. This result was probably already proved, but we could not find a reference, therefore, we added a proof here. A discussion about similar cases may be found in [Noy98].
@@
Let $G$ be $n$-nc tree with vertices $10\leq j\leq n-$ 1.
In the case where $n$ is odd, Lemma 3.7 tells us that there are precisely $n$ elements in every rotation class, so that $nA^{r}(n)=A(n)$.
When $n$ is even, every nc tree that has a rotational symmetry necessarily has $m:=\frac{n}{2}$ elements in its rotation class. Therefore, to compute $A^{r}(n)$, we need to count the number of nc trees that have a rotational symmetry to artificially add them up before dividing by $n$. To count such $G$’s, by Lemma 3.7, there exists $\alpha$ be the rotation of order $2$, and $f=(0,m)$ its fixed edge. Since $f$ splits $G$ in two $(m+1)$-nc trees $G_{1}$ and $G_{2}$ and that those two trees mutually determine the other one via $\alpha$, we’re left to compute the number of $(m+1)$-nc trees $G_{1}$ that have $(0,m)$ as an edge. Using the same argument as in Lemma 3.4 and considering the dual tree tells us that this number is $A_{1}(m+1)$. Therefore, we obtain
[…]
@@ 3.2. Characterization of the topological phase portrait with noncrossing trees
Given a generic polynomial $\overline{P}\in\Xi^{\prime}_{k+1}$, the combinatorial invariant of $\overline{P}{\partial\partial z}$ is the $(k+2)$-nc tree induced by the separatrix graph, with root at $\exp\bigl{(}{{2\pi i2(k+2)}}\bigr{)}$ on the circle at infinity.
@@
Let $\overline{P_{1}},\overline{P_{2}}\in\Xi^{\prime}_{k+1}$. If $\overline{P_{1}}{\partial\partial z}$ and $\overline{P_{2}}{\partial\partial z}$ have the same combinatorial invariant, then they are quasi-conformally equivalent. (In other words, there exists a quasi-conformal mapping $h$ that maps orbits of $\overline{P_{1}}{\partial\partial z}$ on orbits of $\overline{P_{2}}{\partial\partial z}$). In particular, the two vector fields are topologically orbitally equivalent.
@@
$\textstyle{t_{i}^{j}}_{*}\bigl{(}\overline{P_{i}}\frac{\partial}{\partial z}\bigr{)}=\displaystyle|P_{i}|^{2}\frac{\partial}{\partial t_{i}^{j}},$
@@
Recall that the rectifying coordinate $t$ maps orbits to horizontal lines. Since the $\psi_{j}$’s we’ve just built map horizontal lines to horizontal lines, the composition $h_{j}:=(t_{2}^{j})^{-1}\circ\psi_{j}\circ t_{1}^{j}$ maps orbits $\overline{P}_{1}{\partial\partial z}$ in $X_{1}^{j}$ to orbits of $\overline{P}_{2}{\partial\partial z}$ in $X_{2}^{j}$.
Denote by $\mu_{j}(z)=\frac{\overline{\partial}\psi_{j}(z)}{\partial\psi_{j}(z)}$ the Beltrami coefficient of $h_{j}$. One can compute that $\mu_{j}(z)=\frac{\tau_{2}-\tau_{1}}{\overline{\tau_{1}}-\tau_{2}}$ and the condition $|\mu_{j}(z)|<1$ is seen to be equivalent to $04\Im(\tau_{1})\Im(\tau_{2})>$ 0 which is true. Note that all $\psi_{j}$’s are invertible and extend to homeomorphisms between $\overline{A_{1}^{j}}$ and $\overline{A_{2}^{j}}$, where $A_{i}^{j}$ is either a strip or a half-plane. Hence, we have the following diagram for every $j$.
@@
Since $t_{1}^{j}$, $\psi_{j}$ and $t_{2}^{j}$ extend to the separatrices, so does $h_{j}$, so we can glue the $h_{j}$’s in a global homeomorphism $\Psi:\mathbb{C}\to\mathbb{C}$ that maps orbits of $\overline{P_{1}}{\partial\partial z}$ to orbits of $\overline{P_{2}}{\partial\partial z}$. We know that $\Psi$ is quasi-conformal on $\mathbb{C}\setminus\Gamma_{P_{1}}$, hence is quasi-conformal on the whole plane $\mathbb{C}$. ∎
@@
Suppose that $\overline{P}_{j}{{\partial\partial z}}$ has combinatorial invariant $G_{j}$, for $j=1,2$. If there exists a rotation $\rho$ such that $\rho.G_{1}=G_{2}$, then $\rho_{*}\left(\overline{P}_{1}{\partial\partial z}\right)$ is topologically orbitally equivalent to $\overline{P}_{2}{\partial\partial z}$.
As a consequence, there is $A^{r}(k+2)$ generic strata in $\Xi_{k+1}^{\prime}$ modulo rotation.